OpenStudy (anonymous):

2 radical 2 divided by radical 5

OpenStudy (freckles):

\[\frac{2 \sqrt{2}}{\sqrt{5}}\] Do you want to rationalize the denominator ?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

Try multiply the top and bottom by what is on bottom

OpenStudy (freckles):

\[\frac{2 \sqrt{2}}{\sqrt{5}} \cdot \frac{\sqrt{5}}{\sqrt{5}}=\frac{?}{?}\]

OpenStudy (freckles):

what does sqrt(5)*sqrt(5) give us?

OpenStudy (anonymous):

radical 2 over 5

OpenStudy (freckles):

well sqrt(5)*sqrt(5) is 5

OpenStudy (anonymous):

Oh okay, thanks

OpenStudy (freckles):

\[\frac{2 \sqrt{2} \sqrt{5}}{5}\]

OpenStudy (freckles):

so what is sqrt(2)*sqrt(5)?

OpenStudy (freckles):

of what is sqrt(2*5)=?

OpenStudy (freckles):

2*5=?

OpenStudy (anonymous):

10

OpenStudy (freckles):

so sqrt(2)*sqrt(5)=sqrt(10)

OpenStudy (freckles):

in symbols \[\sqrt{2} \cdot \sqrt{5}=\sqrt{10}\]

OpenStudy (anonymous):

oh okay. I kinda get it now..